OpenStudy (anonymous):

Simplify: (x^-2 - x^2)/( x^-1 - x)

OpenStudy (anonymous):

First combine like terms, so X^-2+X^-1=x^-3, Then X^2 +X= x^3, then divide, so the answer should be x^-1